OpenStudy (anonymous):

how do you find an equivalent fraction for 5/6

OpenStudy (sandra):

hey! so to find an equivalent fraction, remember that if you multiple *any* number by 1, you get an equivalent form of the same number - e.g. 2*1 = 2

OpenStudy (sandra):

you can use this logic with fractions too

OpenStudy (sandra):

so for 5/6, I can multiply by 5/5, 6/6, 2/2 - all different versions of 1

OpenStudy (sandra):

and I know I'll have an equivalent fraction: so 5/6 = 5/6*2/2 = 10/12

OpenStudy (sandra):

5/6 = 5/6 * 3/3 = (5*3)/(6*3) = 15/18